1. A U-238 atom (uranium-238) has a nucleus with 92 protons and an atomic mass of 238.0507882 u. What is the nuclear binding energy per nucleon in this atom?

2. When Dr. Bruce Banner becomes angry and undergoes a transformation into the Incredible Hulk, his blood plasma becomes saturated with Na-24 atoms. These radioactive sodium atoms decay into stable Mg-24 by beta decay, along with the release of two high-energy gamma rays, to which Dr. Banner is immune. After initially becoming angry and transforming into the Hulk, how long will it take so that only 15% of the Na-24 atoms remain in his blood plasma? The half-life of Na-24 is 15 hours.
Answer #1

Solution, U-238 A = 238 1 atom pubbliches = Proton = 99 mass - 238 .0507882 atomic neutron = 238-99 = 146 EBIN (ZMH + NMN -